Write an equation of the line that passes through the point and is a) parallel to the given line and b) perpendicular to the line given.
(9,5) (x=8)


I'm not sure about this problem. I don't know whether the line x=8 doesn't have a slope..or is actually the slope itself. If that makes any sense. =P

Found 2 solutions by MRperkins, drk:
Answer by MRperkins(300)

I can see where you are confused. The line x=8 is a vertical line. Because it is vertical, its slope is undefined. Do you remember the vertical line test? Well, this does not pass the vertical line test. Anyways, lets look at what might be parallel to the line. Well, all vertical lines are parallel to each other (all horizontal lines are parallel to all other horizontal lines as well). So we need something that is a line that is just x=? and runs through the given point. Well, how about the line x=9. This is a line that runs through the point (9,5) and is parallel to x=8.
Now for a perpendicular line. Well, vertical lines are perpendicular to horizontal lines. Since we have a vertical line of x=8, we need something perpendicular to it (a horizontal slope) Well, what about y=5? It is perpendicular to the vertical line and it runs through the point (9,5).
.
Do you want your math questions answered quickly and completely? Do you want to ask follow up questions like, “How do you do that?” or “Why did you do that?”? Email your questions to justin.sheppard.tech@hotmail.com and I will answer your questions within 12 hours of submission (usually within 2-3 hours of submission). I charge $2 per question and I will provide a .pdf file of the problem worked out. I will also offer to discuss the problem with you over the phone, through skype, or via an online whiteboard. Payments to be made securely through PayPal AFTER you are satisfied with the answer and services that I provide. This eliminates the possibility of being scammed (worst case scenario is that I provide you with your answer and you don’t pay me).
.
Respectfully,
Justin
justin.sheppard.tech@hotmail.com

Answer by drk(1908)

The line you are give is x = 8. The slope of x = 8 is undefined, or in fractional form 1/0. Now our coordinate is (9,5).
(a) parallel to the given line passing through the point.
step 1 - parallel means same slope, so we can use 1/0.
step 2 - put information into y = mx + b.
5 = (1/0)(9) + b
multiply by side by 0, and you get 5*0 = 1/0*9*0 + b*0 - -> 0 = 9. Now I know we technically can't do that, but I think it helps to see that there will be no y intercept. So the line is
x+=+9.
(b) perpendicular to the given line passing through the point.
step 1 - perpendicular means take the given slope and apply negative flip. S0,
1/0 - - > -0/1 = 0.
step 2 - put information into y = mx + b.
5 = 0*9 + b - - > b = 5. So, we get our answer as
Y+=+5.